Grade 11Mechanics

A stone is thrown vertically upwards. When stone is at half height of its maximum height, it`s speed is 10 ,then maximum height attained by stone is

Shubham
9 Years agoGrade 11

1 Answer

Vaibhav Singh
9 Years ago
This is a tricky one if  you go to solve it by  kinematic formulae.whereas it can be easily solved by energy conservation.
Let maximum height be h therefore half height will be h/2.
  1. Let the initial point be the reference point, therefore potential energy, mgh=0 and kinetic energy KE = 1/2mu
  2. Potential energy at half height, PE = mgh/2 and KE = 1/2m(10)2
  3. PEat max height, PE = mgh and KE = 0.
  4. Now applying energy conservation
PE + KE (at initial) = PE + KE (at half height)
=> 1/2mu2 +0 = 1/2m(100) +mgh/2
=> 1/2u2 = 50 + 5h ( g = 10m/s2)
=> u2 = 100 +10h -------------(i)
Now at heighest point, mgh + 0 = 1/2mu2
=> 10h =1/2u2
=> u2 = 20h -----------(ii)
From (i) and (ii) 
20h = 100 + 10h
=> h = 10 m
Therefore max height is 10 m.